OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3)

hartnn (hartnn):

-2(x-3) distribute -2 what u get ?

OpenStudy (anonymous):

-2x + 6

hartnn (hartnn):

yes, what about 5 – (x + 3)

OpenStudy (anonymous):

5 -x +(-3)??

hartnn (hartnn):

thats 5-x-3 = 2-x right ?

OpenStudy (anonymous):

uhh i don't know :b where u get the 2-x?

hartnn (hartnn):

5-x-3 = (5-3) -x combining like terms

OpenStudy (anonymous):

2 - x oh i see :)

hartnn (hartnn):

yes. so u have now -2x+6 ≥ 2-x can u solve further ?

OpenStudy (anonymous):

ya sure add x to both sides -2x + 6 >= 2 - x +x +x ------------------- -x + 6 >= 2 then aubtract 6 from both sides -6 -6 --------------- -x = -6 Am I right?

hartnn (hartnn):

right side,2-6 = ?

OpenStudy (anonymous):

-4 sorry

hartnn (hartnn):

so -x ≥ -4 then ?

OpenStudy (anonymous):

divide??

OpenStudy (anonymous):

The sign would turn the other way right?

hartnn (hartnn):

divide or multiply by -1 but when u do so, flip the sign.. yes! u know it :)

OpenStudy (anonymous):

wohoo alright thanks hartnn!! so it would become positive x <= 4

hartnn (hartnn):

absolutely!

hartnn (hartnn):

x\(\le\)4

OpenStudy (anonymous):

-2(x - 3) ≥ 5 - (x + 3) -2x + 6 ≥ 5 - x - 3 -x ≥ -4 x ≤ 4